OpenStudy (anonymous):

Can someone help solve these problems? Write the series using sigma notation. -9 + 3 -1 + 1/3 + ...

OpenStudy (anonymous):

Each successive term is reduced by a factor of 3, so the common ratio is 1/3. The signs are also alternating, so you have (-1) to some power. \[\sum_{n=0}^\infty9(-1)^{n+1}\left(\frac{1}{3}\right)^n\\ \sum_{n=0}^\infty-9\left(-\frac{1}{3}\right)^n\]